Horatio Inc. has three divisions which are operated as profit centers. Actual operating data for the divisions listed alphabetically are as follows.

(a)

Compute the missing amounts.

Operating Data

Women’s Shoes

Men’s Shoes

Children’s Shoes

Contribution margin

$312,120

$

(3)

$208,080

Controllable fixed costs

115,600

(4) (5)

Controllable margin

(1)

104,040

109,820

Sales

693,600

520,200

(6)

Variable costs

(2)

369,920

289,000
Answer #1

contribution margin=Sales-Variable cost

contribution margin-Controllable fixed costs=Controllable margin

Women Shoes Men Shoes Children shoes
contribution margin 312120 (520200-369920)=$150,280 208080
Controllable fixed costs 115600 (150280-104040)=$46240 (208080-109820)=$98260
Controllable margin (312120-115600)=$196520 104040 109820
Sales 693600 520200 (208080+289000)=$497,080
Variable costs (693600-312120)=$381480 369920 289,000